Gabriel is 4 years older than Nicole. In 5 years the sum of their ages will be 66. How old is Gabriel now?
Found 2 solutions by ikleyn, Edwin McCravy:
Answer by ikleyn(52803)

.

An equation to start with is


    G + (G-4) = 66 - 5 - 5


    2G - 4    = 56

    2G        = 56 + 4

    2G        = 60

     G        = 60/2 = 30.


ANSWER.  Gabriel is 30 years old now.


Answer by Edwin McCravy(20060)

Ikleyn's is correct, but here is the way I would do it.  I think it best
when two quantities are mentioned, to always choose the SECOND quantity
mentioned to be represented by an unknown letter, not the first one
mentioned, even though that was the one asked for.
Gabriel is 4 years older than Nicole.
 
The SECOND quantity mentioned in that sentence is Nicole's age, so Nicole's
age is the one that I think we should use the unknown N to represent. So

N = Nicole's age now.

Then we define the quantity mentioned first in the sentence, Gabriel's age,
in terms of Nicole's by that sentence
Gabriel is 4 years older than Nicole.
So Gabriel's age is N+4.
In 5 years,...
 
We add 5 to each of their ages:

N+5 = Nicole's age in 5 years.
(N+4)+5 = N+4+5 = N+9 = Gabriel's age in 5 years.
...the sum of their ages will be 66.
(N+5)+(N+9) = 66
    N+5+N+9 = 66
      2N+14 = 66
         2N = 52
          N = 26   <-- Nicole's age now.
Gabriel is 4 years older than Nicole.
So Gabriel is 26+4 = 30.

Edwin